The energy radiated per unit surface area (across all wavelengths) for a black body with temperature...

  1. The energy radiated per unit surface area (across all wavelengths) for a black body with temperature 2200. Use 5.67 x 10-8 for the Stefan-Boltzmann constant.
    1. The Stefan-Boltzmann Law describes the power radiated from a black body in terms of its temperature. Specifically, the total energy radiated per unit surface area of a black body across all wavelengths per unit time is proportional to the fourth power of the black body's thermodynamic temperature
Answer #1

a) energy per unit area = \sigmaT^4

where \sigma = stefan boltzman constant

T = temperature in kelvin

E/A = 5.67*10^-8*(2200)^4

= 1.33*10^6 W/m^2

b) energy per unit area = \sigmaT^4